The difference of two numbers is 91. The second is 5 less than 5 times the first. What are the two numbers?

Answer by ptaylor(1894)


Let x=first number
Then 5x-5=second number
Now we are told that the difference between the two numbers is 91. So:
(5x-5)-x=91 get rid of parens
5x-5-x=91 add 5 to both sides
5x-5+5-x=91+5 collect like terms
4x=96 divide both sides by 4
x=24--------------first number
5x-5=5*24-5=115----------------second number
CK
difference is 91
115-24=91
91=91
Hope this helps----ptaylor